OpenStudy (anonymous):

Solve for x: −2(x + 3) = 4(x − 6)

OpenStudy (anonymous):

use distributive property -2x-6 = 4x-24 add 2x to both sides -6 = 4x -24 add 24 to both sides 6x = 18 divide both sides by 6 x = 3

OpenStudy (anonymous):

Yeah, that's what I got.

OpenStudy (anonymous):

whoops that is -6 = 6x - 24 in the second step

OpenStudy (anonymous):

thankyou!:D